# COSMOLOGICAL DYNAMICS AND OBSERVATIONAL CONSTRAINTS ON A VIABLE $f(Q)$ NON-METRIC GRAVITY MODEL

### 19 Jan 2024 v2

@@ 3 Cosmological dynamics in late-time
In the left panel of Fig. 1 we show the cosmological evolution of $w_{\rm{DE}}$ as a function of $z$ using some fixed positive values for $b$ and we can see that for large values of $z$ and for the distant future ($z\approx-1$), $w_{\rm{DE}}$ is very close to the $\Lambda$CDM model prediction (i.e. $w_{\Lambda}=-1$). For small values of $z$ (the DE era), $w_{\rm{DE}}$ present a variation about -1, and as $b$ decreases, $w_{\rm{DE}}$ progressively converges towards the $\Lambda$CDM model, implying that this variation represents the deviation from $\Lambda$CDM model and it is directly related to the parameter $b$. A similar behavior was found for the $f(R)$ model studied in Ref. 54, but in our case the equation of state does not cross the barrier of -1, i.e. for $b>0$, $w_{\rm{DE}}$ shows a quintessence like behavior.
In the right panel of Fig. 1, the evolution of the same parameter is shown for $b<0$. The same general behavior as for positive $b$ is observed, however $w_{\rm{DE}}$ exhibits now a phantom like behavior, i.e. $w_{\rm{DE}}<-1$ for all values of $z$. A similar behavior for $w_{\rm{DE}}$ was found in Refs. 28 and 30, but in the context of a different exponential $f(Q)$ theories and a numerical solution of the field equations.
@@ 4 Parameter constraints
To constraint the parameters of the model we implement a statistical analysis based in the Markov Chain Monte Carlo (MCMC) method, considering the observational data of the Hubble parameter, $H_{{\rm obs}}(z)$, obtained from the cosmic chronometers and the radial BAO size methods (see Ref. 58 for the complete list of data points and information about the observational methods), and of the magnitude-redshift relation, $m_{{\rm obs}}(z)$, for the Pantheon sample including 1048 observational data [59].
As exposed by Eq. (23), our model depends on two dimensionless parameters ($n,b$); however, we are considering the specific case of $n=1$. Then, for the statistical analysis presented in this section, we also decided to include the current value of the Hubble parameter, $H_{0}\equiv H(z=0)$, as well as the current value of the matter density parameter, $\Omega_{m0}\equiv\Omega_{m}(z=0)$. With this, we compute the value of the Hubble parameter at different redshifts as predicted by the model, $H_{{\rm fQ}}(z)$, as well as the apparent magnitude $m_{{\rm fQ}}(z)$ (which is a function of the expansion rate, $H_{{\rm fQ}}(z)$ through the luminosity distance; see [59] for details), considering the approximate solution given by Eq. (32).
@@
$\mathcal{L}(D|z;\bm{\theta})=-\ln p(D|z;\bm{\theta})=-\frac{1}{2}\chi^{2}(z;\bm{\theta}),$ (41)
@@
$\chi^{2}(z;\bm{\theta})=\sum_{k=1}^{N_{d}}\frac{\left(D_{{\rm fQ},k}(z;\bm{\theta})-D_{{\rm obs},k}(z)\right)^{2}}{\sigma_{k}^{2}},$ (42)
@@
[…]
@@
[…]
@@
$\chi^{2}(z;\bm{\theta})=\sum_{k=1}^{40}\frac{\left(H_{{\rm fQ},k}(z;\bm{\theta})-H_{{\rm obs},k}(z)\right)^{2}}{\sigma(H)_{k}^{2}}+\sum_{k=1}^{1048}\frac{\left(m_{{\rm fQ},k}(z;\bm{\theta})-m_{{\rm obs},k}(z)\right)^{2}}{\sigma(m)_{k}^{2}},$ (43)
in Eq. (41). The results of running emcee in this case are presented in the fourth row of Table 4, and in the corresponding (green) contours and histograms in Fig. 5. As discussed before regarding the prediction for $H_{0}$, in this case the model prefers a value which is closer to the one measured by Planck, and to make a better fit to the data, $\Omega_{m0}$ (within its $1\sigma$ C.L.) is found to be compatible with the $\Lambda$CDM prediction. Also, the perturbation parameter $b$ is found to be negative, excluding $b=0$ at $>2\sigma$ C.L.
@@
[…]
@@ 5 Conclusions
In this sense, here we have introduce a new and viable $f(Q)$ gravity model which can be represented as a perturbation of the $\Lambda$CDM model (see Eq. (23)). Usually, within the realm of $f(Q)$ gravity, the customary approach to investigate cosmological evolution involves employing a parametrization of the Hubble parameter in terms of redshift, among other strategies. In this work as an alternative, we have derived an analytical approximation for the expansion rate $H(z)$ (Eq. (32)), and from it, we have deduced approximated analytical expressions for the parameters $w_{\rm{DE}}$, $w_{\rm{eff}}$, and $\Omega_{DE}$, as well as the deceleration parameter $q$ in terms of the redshift $z$ (Eqs. (37)-(40)).
@@
Acknowledgements
A. O. is supported by Patrimonio Autónomo-Fondo Nacional de Financiamiento para la Ciencia, la Tecnología y la Innovación Francisco José de Caldas (MINCIENCIAS-COLOMBIA) Grant No. 110685269447 RC-80740-465-2020, projects 69723 and 69553. We also thank Ricardo Vega (Universidad del Atlántico) for helping us with the computer resources of his Laboratory used for the MCMC analyses.

### 30 Jan 2024 v3

@@ 3 Cosmological dynamics in late-time
In the left panel of Fig. 1 we show the cosmological evolution of $w_{\rm{DE}}$ as a function of $z$ using some fixed positive values for $b$ and we can see that for large values of $z$ and for the distant future ($z\approx-1$), $w_{\rm{DE}}$ is very close to the $\Lambda$CDM model prediction (i.e. $w_{\Lambda}=-1$). For small values of $z$ (the DE era), $w_{\rm{DE}}$ present a variation about -1, and as $b$ decreases, $w_{\rm{DE}}$ progressively converges towards the $\Lambda$CDM model, implying that this variation represents the deviation from $\Lambda$CDM model and it is directly related to the parameter $b$. A similar behavior was found for the $f(R)$ model studied in Ref. [54], but in our case the equation of state does not cross the barrier of -1, i.e. for $b>0$, $w_{\rm{DE}}$ shows a quintessence like behavior.
In the right panel of Fig. 1, the evolution of the same parameter is shown for $b<0$. The same general behavior as for positive $b$ is observed, however $w_{\rm{DE}}$ exhibits now a phantom like behavior, i.e. $w_{\rm{DE}}<-1$ for all values of $z$. A similar behavior for $w_{\rm{DE}}$ was found in Refs. [28] and [30], but in the context of a different exponential $f(Q)$ theories and a numerical solution of the field equations.
@@ 4 Parameter constraints
To constraint the parameters of the model we implement a statistical analysis based in the Markov Chain Monte Carlo (MCMC) method, considering the observational data of the Hubble parameter, $H_{{\rm obs}}(z)$, obtained from the cosmic chronometers and the radial BAO size methods (see Ref. [58] for the complete list of data points and information about the observational methods), and of the magnitude-redshift relation, $m_{{\rm obs}}(z)$, for the Pantheon sample including 1048 observational data [59].
As exposed by Eq. (3.1), our model depends on two dimensionless parameters ($n,b$); however, we are considering the specific case of $n=1$. Then, for the statistical analysis presented in this section, we also decided to include the current value of the Hubble parameter, $H_{0}\equiv H(z=0)$, as well as the current value of the matter density parameter, $\Omega_{m0}\equiv\Omega_{m}(z=0)$. With this, we compute the value of the Hubble parameter at different redshifts as predicted by the model, $H_{{\rm fQ}}(z)$, as well as the apparent magnitude $m_{{\rm fQ}}(z)$ (which is a function of the expansion rate, $H_{{\rm fQ}}(z)$ through the luminosity distance; see [59] for details), considering the approximate solution given by Eq. (3.10).
@@
$\mathcal{L}(D|z;\bm{\theta})=-\ln p(D|z;\bm{\theta})=-\frac{1}{2}\chi^{2}(z;\bm{\theta}),$ (4.1)
@@
$\chi^{2}(z;\bm{\theta})=\sum_{k=1}^{N_{d}}\frac{\left(D_{{\rm fQ},k}(z;\bm{\theta})-D_{{\rm obs},k}(z)\right)^{2}}{\sigma_{k}^{2}},$ (4.2)
@@
Table 1: Intervals used as the (flat) priors for the parameter studied in this analysis, $\left(H_{0},\Omega_{m0},b\right)$.
$\theta_{i}$ $H_{0}$ (km/s/Mpc) $\Omega_{m0}$ $b$
Prior (60, 80) (0.1, 0.4) (-2, 2)
@@
Table 2: Resulting $1\sigma$ allowed intervals for the fitted parameters from the MCMC analysis. The values for the $\Lambda$CDM model are from Planck 2018 [61].
Model $H_{0}$ (km/s/Mpc) $\Omega_{m0}$ $b$
$\Lambda$CDM $67.36\pm 0.54$ $0.315\pm 0.007$
$f(Q):H(z)$ $67.79_{-5.35}^{+7.56}$ $0.279_{-0.060}^{+0.059}$ $-0.116_{-0.362}^{+0.309}$
$f(Q):$ Pantheon $68.25_{-4.80}^{+4.22}$ $0.238_{-0.038}^{+0.135}$ $0.163_{-0.430}^{+0.425}$
$f(Q):H(z)+$Pantheon $64.41_{-0.61}^{+0.67}$ $0.330\pm 0.018$ $-0.151_{-0.069}^{+0.068}$
@@
$\chi^{2}(z;\bm{\theta})=\sum_{k=1}^{40}\frac{\left(H_{{\rm fQ},k}(z;\bm{\theta})-H_{{\rm obs},k}(z)\right)^{2}}{\sigma(H)_{k}^{2}}+\sum_{k=1}^{1048}\frac{\left(m_{{\rm fQ},k}(z;\bm{\theta})-m_{{\rm obs},k}(z)\right)^{2}}{\sigma(m)_{k}^{2}},$ (4.3)
in Eq. (4.1). The results of running emcee in this case are presented in the fourth row of Table 2, and in the corresponding (green) contours and histograms in Fig. 5. As discussed before regarding the prediction for $H_{0}$, in this case the model prefers a value which is closer to the one measured by Planck, and to make a better fit to the data, $\Omega_{m0}$ (within its $1\sigma$ C.L.) is found to be compatible with the $\Lambda$CDM prediction. Also, the perturbation parameter $b$ is found to be negative, excluding $b=0$ at $>2\sigma$ C.L.
@@
Table 3: Present ($z=0$) value of $w_{\rm{DE}}$ and $\Omega_{\rm{DE}}$ as obtained from the proposed model, compared with the Planck 2018 results [61].
Model $w_{\rm{DE}0}$ $\Omega_{\rm{DE}0}$
$\Lambda$CDM $-1.03\pm 0.03$ $0.6847\pm 0.0073$
$f(Q):H(z)$ $-1.02$ $0.7458$
$f(Q):$ Pantheon $-0.93$ $0.7280$
$f(Q):H(z)+$Pantheon $-1.03$ $0.7029$
@@ 5 Conclusions
In this sense, here we have introduce a new and viable $f(Q)$ gravity model which can be represented as a perturbation of the $\Lambda$CDM model (see Eq. (3.1)). Usually, within the realm of $f(Q)$ gravity, the customary approach to investigate cosmological evolution involves employing a parametrization of the Hubble parameter in terms of redshift, among other strategies. In this work as an alternative, we have derived an analytical approximation for the expansion rate $H(z)$ (Eq. (3.10)), and from it, we have deduced approximated analytical expressions for the parameters $w_{\rm{DE}}$, $w_{\rm{eff}}$, and $\Omega_{DE}$, as well as the deceleration parameter $q$ in terms of the redshift $z$ (Eqs. (3.15)-(3.18)).
@@
Acknowledgements A. O. is supported by Patrimonio Autónomo-Fondo Nacional de Financiamiento para la Ciencia, la Tecnología y la Innovación Francisco José de Caldas (MINCIENCIAS-COLOMBIA) Grant No. 110685269447 RC-80740-465-2020, projects 69723 and 69553. We also thank Ricardo Vega (Universidad del Atlántico) for helping us with the computer resources of his Laboratory used for the MCMC analyses.